# Sample Complexity of Diffusion Model Training Without Empirical Risk Minimizer Access

### v2

@@ 3 Sample Complexity of Diffusion Models
The data distribution $p_{0}$ of the data variable $x_{0}$ has an absolutely continuous CDF, is supported on a continuous set $\Gamma\in\mathbb{R}^{d}$, and there exists a constant $0<C_{1}<\infty$ such that for every $t\geq 0$ we have $P(|x_{0}|\geq t)\leq 2\cdot{\exp}^{-\frac{t^{2}}{C^{2}_{1}}}$.
[10, 32] assume that the data distribution is supported on a bounded set, thereby excluding commonly encountered distributions such as Gaussian and sub-Gaussian families. In contrast, our analysis only requires the data distribution to be sub-Gaussian, making our results applicable to a significantly broader class of distributions.
Together, these assumptions form a minimal yet sufficient foundation for analyzing score estimation in practice. Smoothness and bounded gradient variance implied by the sub-Gaussian assumption are mild and generally satisfied for standard neural architectures with ReLU or GELU activations and well-behaved data distributions. The PL condition has been shown to emerge in over-parameterized networks or under lazy training regimes, where the function class is expressive enough to approximate the ground-truth score function [28]. Notably, these conditions are not only specific to our setting they have been widely adopted in recent works studying the optimization landscape of deep diffusion models [34, 29]. Under these assumptions, we now characterize the sample complexity of score-based learning in the following.
Main Result: To estimate the score function $s_{t}(x)$, modern diffusion models train a neural network to predict the noise $\epsilon_{\theta}(x_{t},t)$ injected in the forward process. The training objective, as used in DDPM [16], corresponds to minimizing
$\mathcal{L}(\theta)=\mathbb{E}_{x,t,\epsilon}\left[\|\epsilon-\epsilon_{\theta}(x_{t},t)\|^{2}\right],$ (8)
where $x_{t}=\sqrt{\bar{\alpha}_{t}}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon$ is the noisy observation at time $t$, and $\bar{\alpha}_{t}=\prod_{s=1}^{t}(1-\beta_{s})$ is defined by the noise schedule $\{\beta_{t}\}_{t=1}^{T}$. This objective aligns with the score-matching loss under a reparameterized formulation. Algorithm 1 (see Appendix C) outlines the standard DDPM training and sampling pipeline. The final generative quality depends on the fidelity of the learned score estimate across time steps.
The following theorem provides an upper bound on the total variation distance between the true data distribution $p$ and the generated distribution $\hat{p}_{t_{0}}$.
Theorem 1 (Total Variation Distance Bound).
Let $p(x)$ denote the true data distribution, and $\hat{p}_{t_{0}}(x)$ be the distribution generated by a discretized diffusion model with time steps $\{t_{k}\}_{k=0}^{K}$, reverse-time score estimates $\hat{s}_{T-t_{k}}(x)$, and stopping at time $t_{0}$. Let $d$ be the data dimension, and $n_{k}$ be the number of samples for score estimation at time step $t_{k}$.
Assume the loss function $\mathcal{L}_{t}(\theta)$ satisfies Assumptions 1 and 2, the data distribution satisfies Assumption 3, and the learning rate satisfies $0\leq\eta\leq\frac{1}{\kappa}$. Further assume
$n_{k}=\mathcal{O}\left(d^{2}.\log\left(\frac{4K}{\delta}\right)\left(\frac{\epsilon^{-6}}{\sigma_{k}^{-4}}\right)\right),$ (9)
and that the neural network has width $W$ and depth $D$ such that $WD=\mathcal{O}\left((\epsilon/\sigma_{k})^{-\frac{d}{6}}\right)$.
Then, with probability at least $1-\delta$, the total variation distance between the true distribution $p$ and $\hat{p}_{t_{0}}$ satisfies
$\mathrm{TV}(p,\hat{p}_{t_{0}})\leq\sqrt{t_{0}}\log\left(\frac{1}{t_{0}}\right)+\mathcal{O}(\exp^{-T})+\mathcal{O}\left(\frac{1}{\sqrt{K}}\right)+\left(\epsilon\cdot{T}\right)+\epsilon$ (10)
Furthermore, by setting $T=\mathcal{O}\left(\log\left(\frac{1}{\epsilon}\right)\right)$, $t_{0}=\tilde{\mathcal{O}}(\epsilon^{2})$ and $K=\tilde{\mathcal{O}}(\epsilon^{-2})$, we obtain
$\mathrm{TV}(p,\hat{p}_{t_{0}})\leq\tilde{\mathcal{O}}(\epsilon),$ (11)
@@
Theorem 1 establishes that the total variation distance between the true data distribution and the diffusion model’s output can be made arbitrarily small specifically, $\tilde{\mathcal{O}}(\epsilon)$ by properly scaling model capacity and algorithmic parameters. To the best of our knowledge, these are the only known sample complexity bounds for score-based diffusion models, improving upon the prior results as discussed in the introduction without assuming access to empirical risk minimizer for the score estimation loss. Our result highlight an inherent trade-off between the sample complexity bounds and the assumption of access to the empirical risk minimizer.
@@
Proof of Theorem 1. Recall that $\hat{p}_{t_{0}}$ is derived via score-based sampling, so using the triangle inequality repeatedly to decompose the TV distance between the true distribution $p$ and $\hat{p}_{t_{0}}$
Using the triangle property of $\mathrm{TV}$ we write
$\mathrm{TV}(p,\hat{p}_{t_{0}})$ $\leq\mathrm{TV}(p,p_{t_{0}})+\mathrm{TV}(p_{t_{0}},\hat{p}_{t_{0}}).$ (12)
$\leq\sqrt{t_{0}}\log\left(\frac{1}{t_{0}}\right)+\mathrm{TV}(p_{t_{0}},\hat{p}_{t_{0}}).$ (13)
The bound on $\mathrm{TV}(p,p_{t_{0}})$ follows from Lemma D.5 of [13].
Now lemma B.2 of [15] states that if $P\left(\sum_{k=0}^{K}||\hat{s}_{t_{k}}(x,t_{k})-{\nabla}{\log}p_{t_{k}}(x)||^{2}.(t_{k+1}-t_{k})\geq\epsilon^{2}\right)\leq\beta$, where $\hat{s}_{t_{k}}(x,t_{k})$ is the estimate of the score ${\nabla}{\log}p_{t_{k}}(x)$, then we have
$\mathrm{TV}(p_{t_{0}},\hat{p}_{t_{0}})\leq\beta+\epsilon+\mathrm{TV}(p_{t_{0}},p^{dis}_{t_{0}})+\mathrm{TV}(\widetilde{p}_{t_{0}},p^{\mathrm{dis}}_{t_{0}})$ (14)
Let ${\beta(n)}$ be such that
$\sum_{k=t_{o}}^{K}\mathbb{E}_{x\sim p_{t_{k}}}||\hat{s}_{t_{k}}(x,t_{k})-{\nabla}{\log}p_{t_{k}}(x)||^{2}(t_{k}-t_{k-1})\leq{\beta(n)}$ (15)
From Theorem 2 from [31] we have that
$P(\sum_{k=t_{o}}^{K}||\hat{s}_{t_{k}}(x,t_{k})-{\nabla}{\log}p_{t_{k}}(x)||^{2}\geq\epsilon^{2})\leq\frac{1}{\epsilon^{2}}\cdot\beta(n)$ (16)
Here $n$ denotes the number of samples used to estimate the loss functions $\hat{s}_{t_{k}}(x,t_{k})$ Thus, we obtain that
$\mathrm{TV}(p_{t_{0}},\hat{p}_{t_{0}})\leq\delta+\epsilon+\mathrm{TV}(p_{t_{0}},p^{dis}_{t_{0}})+\mathrm{TV}(\widetilde{p}_{t_{0}},p^{\mathrm{dis}}_{t_{0}})$ (17)
$\leq\beta(n).\frac{1}{\epsilon}+\epsilon+\mathrm{TV}(p_{t_{0}},p^{dis}_{t_{0}})+\mathrm{TV}(\widetilde{p}_{t_{0}},p^{\mathrm{dis}}_{t_{0}})$ (18)
@@
The bounds on $\mathrm{TV}(p^{\prime}_{t_{0}},p^{\mathrm{dis}}_{t_{0}})$ and $\mathrm{TV}(p_{t_{0}},p^{\prime}_{t_{0}})$ follow from Lemma B.4 of [15] and Proposition 4 of [5], respectively to get
$\mathrm{TV}(p,\hat{p}_{t_{0}})$ $\leq\sqrt{t_{0}}\log\left(\frac{1}{t_{0}}\right)+\exp(-T)+{\mathcal{O}}\left(\frac{1}{\sqrt{K}}\right)+\beta(n)\cdot\frac{1}{\epsilon^{2}}+\epsilon$ (19)
@@
Note that we have used results from [15], [5] and [13]. [15] and [5] assume a bounded second moment for the data distribution and [13] assumes sub-Gaussian structure for the data distribution. Both of these assumptions are together given in Assumption 3. Putting the expressions for $\mathrm{TV}(p^{\prime}_{t_{0}},p^{\mathrm{dis}}_{t_{0}})$, and $\mathrm{TV}(p_{t_{0}},p^{\prime}_{t_{0}})$ in Equation (18), we get
A key point here is that in [15], the term $\sum_{k=0}^{K}\|\hat{s}_{t_{k}}(x,t_{k})-\nabla\log p_{t_{k}}(x)\|^{2}(t_{k+1}-t_{k})$ was required to be bounded by $\epsilon^{2}$ with high probability in order for the upper bound on the total variation distance between the learned and true distributions to hold. In contrast, we relax this requirement by converting the probabilistic bound into a bound on the expectation of the same quantity. Specifically, we show that it suffices for the expectation of $\sum_{k=0}^{K}\|\hat{s}_{t_{k}}(x,t_{k})-\nabla\log p_{t_{k}}(x)\|^{2}(t_{k+1}-t_{k})$ to be upper bounded by $\beta(n)$ (See Remark 1 for detail). This shift from high-probability bounds to an expectation-based analysis is a crucial aspect of our approach. In the following, we analyze how $\beta(n)$ depends on the number of samples used to estimate the score function $\hat{s}_{t_{k}}(x)$.
In order to obtain an expression for $\beta(n)$, we denote $A(k)$ as
$A(k)\coloneqq E_{x\sim p_{t_{k}}}\left\|\hat{s}_{t_{k}}(x,t_{k})-\nabla\log p_{t_{k}}(x)\right\|^{2}dx$ (20)
@@
and denote $s^{a}_{k}$ and $s^{b}_{k}$ as the score functions associated with the parameters $\theta^{a}_{k}$ and $\theta^{b}_{k}$ respectively. Approximation error $\mathcal{E}^{\mathrm{approx}}_{k}$ captures the error due to the limited expressiveness of the function class $\{s_{\theta}\}_{\theta\in\Theta}$. The statistical error $\mathcal{E}^{\mathrm{stat}}_{k}$ is the error from using a finite sample size. Finally, the optimization error $\mathcal{E}^{\mathrm{opt}}_{k}$ is due to not reaching the global minimum during training.
One of our key contributions lies in rigorously bounding each of these error components and showing how their interplay governs the overall generative error. In particular, we derive novel bounds that explicitly capture the dependencies on sample size, neural network capacity, and optimization parameters, without any assumption on the access to empirical risk minimizer of the score estimation loss. We formalize these results in the following lemmas. Detailed proofs are deferred to Appendices D.1, D.2, and D.3, respectively.
@@
Let $W$ and $D$ denote the width and depth of the neural network architecture, respectively, and let $d$ represent the input data dimension. Then, under Assumption 3 for all $k\in[{0},K]$, we have
$\mathcal{E}^{\mathrm{approx}}_{k}\leq\mathcal{O}\left(\left(WD\right)^{-\frac{2}{d}}\right).$ (24)
For this lemma we use the conditional normality of the score function conditioned on the initial data observation to account for the fact that the score function can be unbounded. We combine that with approximation upper bounds from [36, 41] to obtain our result. The specific form of decomposition of the score estimation error allows this term to be separate from the error terms depending on the number of samples used for score estimation. This term also accounts for the error due to the limited approximation power of a class of neural networks, an important factor that was not considered in previous analysis.
@@
Let $n_{k}$ denote the number of samples used to estimate the score function at each diffusion step $t_{k}$. Then, under Assumptions 3, for all $k\in[{0},K]$, with probability at least $1-\delta$, we have
$\mathcal{E}^{\mathrm{stat}}_{k}\leq\mathcal{O}\left(d.\sqrt{\frac{{{\log}\left(\frac{2}{\delta}\right)}}{{n_{k}}}}\right)$ (25)
The proof of this lemma follows from utilizing the definitions of $s^{a}_{t}$ and $s^{b}_{t}$. Our novel analysis here allows us to avoid the exponential dependence on the neural networks parameters in the sample complexity. This is achieved without resorting to the quantile formulation and assuming access to the empirical risk minimizer of the score estimation loss which leads to a sample complexity of $\widetilde{\mathcal{O}}(\epsilon^{-5})$ in [15].
@@
Let $n_{k}$ be the number of samples used to estimate the score function at diffusion step ${t_{k}}$. If the learning rate $0\leq\eta\leq\frac{1}{\kappa}$, then under Assumptions 1, 2 and 3, for all $k\in[0,K]$, the optimization error due to imperfect minimization of the training loss satisfies with probability at least $1-\delta$
$\mathcal{E}^{\mathrm{opt}}_{k}\leq\mathcal{O}\left(d.\sqrt{\frac{\log\left(\frac{2}{\delta}\right)}{n_{k}}}\right).$ (26)
We leverage assumptions 1, 2, and 3 alongside our unique analysis to derive an upper bound on $\mathcal{E}_{\mathrm{opt}}$. The key here is our recursive analysis of the error at each stochastic gradient descent (SGD) step, which captures the error introduced by the finite number of SGD steps in estimating the score function. This is the first analysis of diffusion models to explicitly account for this error, whereas prior works assumed no such error, treating the empirical loss minimizer as if it were known exactly.
Combining the decomposition in Equation (21) along with Lemmas 13, we obtain the following bound on $A(k)$ (Equation (20)) with probability at least $1-\delta$
$A(k)$ $\leq\mathcal{O}\left(d.\sqrt{\frac{{{\log}\left(\frac{2}{\delta}\right)}}{{n_{k}}}}\right)+\mathcal{O}\left(d.\sqrt{\frac{\log\left(\frac{2}{\delta}\right)}{n_{k}}}\right)+\mathcal{O}\left((WD)^{-\frac{2}{d}}\right)$
$\leq\mathcal{O}\left(d.\sqrt{\frac{{{\log}\left(\frac{2}{\delta}\right)}}{{n_{k}}}}\right)+\mathcal{O}\left((WD)^{-\frac{2}{d}}\right),$ (27)
where in the second inequality we combine the first two term appropriately. Setting the network size $WD=\mathcal{O}\left((\epsilon/\sigma_{k})^{-\frac{d}{6}}\right)$ and the sample size
$n_{k}=\mathcal{O}\left(d^{2}.\log\left(\frac{4K}{\delta}\right)\left(\frac{\epsilon^{-6}}{\sigma_{k}^{-4}}\right)\right),$ (28)
we ensure that $A(k)\leq\frac{\epsilon^{3}}{\sigma_{k}^{2}}=\frac{\epsilon^{3}}{1-e^{-2(T-t_{k})}}$ for all $k\in\{0,\dots,K-1\}$. Summing over all time steps, we obtain with probability at least $1-\delta$
[…]
Note that the term $\left(\log^{2}\left(\frac{4K}{\delta}\right)\right)$ appears in the upper bound for $n_{k}$ in Equation (28) since we have to take a union bound for Lemma 2 and Lemma 3 and then take a union bound over $K$ discretization steps. Substituting this bound into Equation (20), and subsequently into the decomposition in Equation (14) we get that $\beta(n)\leq\epsilon^{3}T$. Substituting this bound in Equation (19) we obtain that with probability at least $1-\delta$.
$\mathrm{TV}(p,\hat{p}_{t_{0}})$ $\leq\sqrt{t_{0}}\log\left(\frac{1}{t_{0}}\right)+\mathcal{O}(\exp^{-T})+\mathcal{O}\left(\frac{1}{\sqrt{K}}\right)+\epsilon\cdot{T}+{\epsilon}$ (30)
Finally, by choosing $T=\mathcal{O}\left(\log\left(\frac{1}{\epsilon}\right)\right)$, $t_{0}=\tilde{\mathcal{O}}(\epsilon^{2})$ and $K=\tilde{\mathcal{O}}(\epsilon^{-2})$, we conclude that with probability at least $1-\delta$
$\mathrm{TV}(p,\hat{p}_{t_{0}})\leq\tilde{\mathcal{O}}(\epsilon),$ (31)
@@
Remark 1.
The key distinction between our approach and [15] lies in the error aggregation strategy. In [15], each term $\|\hat{s}_{t_{k}}(x)-\nabla\log p_{t_{k}}(x)\|^{2}$ in the summation (Equation (15)) must satisfy $\|\hat{s}_{t_{k}}(x)-\nabla\log p_{t_{k}}(x)\|^{2}\leq\epsilon^{2}$ with probability at least $1-\frac{\beta}{K}$ for all $k\in[0,K]$, necessitating a union bound over $K$ terms (see Lemma C.1 in [15]). This introduces an extraneous factor of $K$ in their sample complexity (evident in Theorem C.2 of [15], detailed in Appendix B). In contrast, we bound the expected cumulative error $\sum_{k=0}^{K}\mathbb{E}_{x\sim p_{t_{k}}}\|\hat{s}_{t_{k}}(x)-\nabla\log p_{t_{k}}(x)\|^{2}(t_{k+1}-t_{k})$.
Error Decomposition: The expectation-centric analysis enables principled decomposition of the score estimation error into approximation, statistical, and optimization components (Equation (21)). This framework explicitly quantifies errors arising from imperfect optimization, a feature absent in previous works, including [15], which assume access to exact empirical risk minimizers.
In summary, our work provides a principled decomposition of the errors in score-based generative models, highlighting how each component contributes to the overall sample complexity. This leads to the first finite sample complexity bound of $\widetilde{\mathcal{O}}(\epsilon^{-6})$ for diffusion models without assuming access to the empirical minimizer of the score estimation function. The sample complexity bound also avoids the exponential dependence on the data dimension or the neural network parameters.

### v5

@@ 3 Sample Complexity of Diffusion Models
The data distribution $p_{0}$ of the data variable $x_{0}$ has an absolutely continuous CDF, is supported on a continuous set $\Gamma\in\mathbb{R}^{d}$, and there exists a constant $0<C_{2}<\infty$ such that for every $t\geq 0$ we have $P(|x_{0}|\geq t)\leq 2\cdot{\exp}^{-\frac{t^{2}}{C^{2}_{2}}}$.
Theorem 2 (Total Variation Distance Bound Under Sub-Gaussian Assumption).
Assume that the data distribution satisfies Assumption 5, the loss function $\mathcal{L}_{k}(\theta)$ satisfies Assumptions 2 3,4 for all $k\in[0,K]$ and the learning rate for the $i^{th}$ SDG step with one sample per step, satisfies $\eta_{i}=\frac{\alpha}{i+\gamma}$ where $\alpha.\mu>1$ and $\gamma>\alpha.\kappa$. Further assume
$n_{k}={\Omega}\left(W^{2D}{\cdot}d^{2}.\log\left(\frac{4K}{\delta}\right)\left(\frac{\epsilon^{-4}}{\sigma_{k}^{-4}}\right)\right),$ (12)
Then, with probability at least $1-\delta$, the total variation distance between the $p_{0}$ and $\hat{p}_{t_{0}}$ satisfies
$TV(p_{0},\hat{p}_{t_{0}})$ $\leq\mathcal{O}\left(\sqrt{t_{0}}\log\!\bigl(1/t_{0}\bigr)\right)+\mathcal{O}(\exp^{-T})+\mathcal{O}\left(\frac{1}{\sqrt{K}}\right)+\mathcal{O}(\epsilon)+\epsilon_{approx}$ (13)
Furthermore, by setting $t_{0}=\mathcal{O}(\epsilon^{-2})$, $T=\Omega\left(\log\left(\frac{1}{\epsilon}\right)\right)$ and $K=\Omega(\epsilon^{-2})$, we obtain
$\mathrm{TV}(p_{0},\hat{p}_{t_{0}})\leq{\mathcal{O}}(\epsilon)+\epsilon_{approx},$ (14)
@@
Proof of Theorem 1.
@@
Recall that $\hat{p}_{t_{0}}$ is derived via score-based sampling, so using the triangle inequality repeatedly to decompose the TV distance between the true distribution $p_{t_{0}}$ and $\hat{p}_{t_{0}}$ as follows
$\mathrm{TV}(p_{t_{0}},\hat{p}_{t_{0}})\leq\mathrm{TV}(p_{t_{0}},p^{dis}_{t_{0}})+\mathrm{TV}(p^{\mathrm{dis}}_{t_{0}},\widetilde{p}_{t_{0}})+\mathrm{TV}(\widetilde{p}_{t_{0}},\hat{p}_{t_{0}},)$ (15)
@@
The bounds on $\mathrm{TV}(p_{t_{0}},p^{dis}_{t_{0}})$ and $\mathrm{TV}(p^{\mathrm{dis}}_{t_{0}},\widetilde{p}_{t_{0}})$ follow from Lemma B.4 of Gupta et al. (2024) and Proposition 4 of Benton et al. (2024), respectively to get
$\mathrm{TV}((p_{t_{0}},\hat{p}_{t_{0}})$ $\leq\exp(-T)+\mathrm{TV}(p^{\mathrm{dis}}_{t_{0}},\widetilde{p}_{t_{0}})+{\mathcal{O}}\left(\frac{1}{\sqrt{K}}\right)$ (16)
@@
Note that we have used results from Gupta et al. (2024) and Benton et al. (2024) which assume a bounded second moment for the data distribution. This is satisfied by Assumption 1. Now from lemma 4, $\mathrm{TV}(p^{\mathrm{dis}}_{t_{0}},\widetilde{p}_{t_{0}})$ is upper bounded as follows
$\mathrm{TV}(p^{\mathrm{dis}}_{t_{0}},\widetilde{p}_{t_{0}})\leq\frac{1}{2}\sqrt{\sum_{k=0}^{K}E_{x\sim p_{t_{k}}}\left\|\hat{s}_{t_{k}}(x,t_{k})-\nabla\log p_{t_{k}}(x)\right\|^{2}(t_{k+1}-t_{k})}$ (17)
In order to upper bound, $\mathrm{TV}(p^{\mathrm{dis}}_{t_{0}},\widetilde{p}_{t_{0}})$ we denote $A(k)$ as
$A(k):=E_{x\sim p_{t_{k}}}\left\|\hat{s}_{t_{k}}(x,t_{k})-\nabla\log p_{t_{k}}(x)\right\|^{2}dx$ (18)
Therefore, bounding the TV distance between $p$ and $\hat{p}_{t_{0}}$ translates to bounding the cumulative error in estimating the score function at different time steps. We now focus on bounding this term. Specifically, we have that
$\mathrm{TV}(p_{t_{0}},\hat{p}_{t_{0}})$ $\leq\exp(-T)+\frac{1}{2}\sqrt{\sum_{k=0}^{K}A_{k}{\cdot}(t_{k+1}-t_{k})dx}+{\mathcal{O}}\left(\frac{1}{\sqrt{K}}\right)$ (19)
@@
and denote $s^{a}_{t_{k}}$ and $s^{b}_{t_{k}}$ as the score functions associated with the parameters $\theta^{a}_{k}$ and $\theta^{b}_{k}$ respectively. Approximation error $\mathcal{E}^{\mathrm{approx}}_{k}$ captures the error due to the limited expressiveness of the function class $\{s_{\theta}\}_{\theta\in\Theta}$. The statistical error $\mathcal{E}^{\mathrm{stat}}_{k}$ is the error from using a finite sample size. Finally, the optimization error $\mathcal{E}^{\mathrm{opt}}_{k}$ is due to not reaching the global minimum during training.
One of our key contributions lies in rigorously bounding each of these error components and showing how their interplay governs the overall generative error. In particular, we derive novel bounds that explicitly capture the dependencies on sample size, neural network capacity, and optimization parameters, without any assumption on the access to the empirical risk minimizer of the score estimation loss. We formalize these results in the following lemmas. Detailed proofs are deferred to Appendices C.1, and C.2, respectively.
@@
$\mathcal{E}_{k}^{\mathrm{approx}}$ is defined as follows
$\mathcal{E}_{k}^{\mathrm{approx}}=\text{min}_{\theta\in\Theta}\mathbb{E}_{x\sim p_{t_{k}}}\left[\|s_{\theta}(x,t_{k})-\nabla\log p_{t}(x,t_{k})\|^{2}\right]$ (23)
Then, under Assumption 3 for all $k\in[{0},K]$, we have
$\mathcal{E}^{\mathrm{approx}}_{k}\leq\epsilon_{approx}$ (24)
This result directly follows from Assumption 3 and the definition of $\mathcal{E}^{\mathrm{approx}}_{k}$.
@@
Let $n_{k}$ denote the number of samples used to estimate the score function at each diffusion step $t_{k}$. If the data distribution satisfies the Assumption 1 and the loss function $\mathcal{L}_{k}(\theta)$ satisfies Assumptions 2 for all $k\in[{0},K]$, then with probability at least $1-\delta$, we have
$\mathcal{E}^{\mathrm{stat}}_{k}\leq{\mathcal{O}}\left(W^{D}.d.\sqrt{\frac{{{\log}\left(\frac{2}{\delta}\right)}}{{n_{k}}}}\right)$ (25)
The proof of this lemma follows from utilizing the definitions of $s^{a}_{t}$ and $s^{b}_{t}$. Existing analyses of statistical errors, such as those given in Shalev-Shwartz & Ben-David (2014), only work when the loss function is bounded. This is not the case for diffusion models. Thus, we use a novel analysis that uses the conditional normality of the score function as well as the bounded second moment property of the data variable in Assumption 1 to obtain the upper bound on the statistical error. The details of the analysis are given in Appendix C.1.
@@
Let $n_{k}$ be the number of samples used to estimate the score function at diffusion step $t_{k}$}. Assume that the score loss function $\mathcal{L}_{k}(\theta)$ satisfies the Assumptions 2 and 4, for all $k\in[0,K]$, and the learning rate for the $i^{th}$ SDG step with one sample per step, satisfies $\eta_{i}=\frac{\alpha}{i+\gamma}$ where $\alpha.\mu>1$ and $\gamma>\alpha.\kappa$, then with probability at least $1-\delta$
$\mathcal{E}^{\mathrm{opt}}_{k}\leq{\mathcal{O}}\left(W^{D}.d.\sqrt{\frac{\log\left(\frac{2}{\delta}\right)}{n_{k}}}\right).$ (26)
We leverage assumptions 2, 4 , alongside our unique recursive at each stochastic gradient descent (SGD) step, which captures the error introduced by the finite number of SGD steps in estimating the score function. This is the first analysis of diffusion models to explicitly account for this error. All prior works assumed no such error, treating the empirical loss minimizer as if it were known exactly. The details of the analysis are given in Appendix C.2.
Combining the decomposition in equation 20 along with Lemmas 13, we obtain the following bound on $A(k)$ ( equation 18) with probability at least $1-\delta$
$A(k)$ $\leq\mathcal{O}\left(W^{D}.d.\sqrt{\frac{{{\log}\left(\frac{2}{\delta}\right)}}{{n_{k}}}}\right)+\mathcal{O}\left(W^{D}.d.\sqrt{\frac{\log\left(\frac{2}{\delta}\right)}{n_{k}}}\right)+\epsilon_{approx}$ (27)
$\leq\mathcal{O}\left(W^{D}.d.\sqrt{\frac{{{\log}\left(\frac{2}{\delta}\right)}}{{n_{k}}}}\right)+\epsilon_{approx},$ (28)
where in the second inequality we combine the first two terms appropriately. Setting the sample size
$n_{k}={\Omega}\left(W^{2D}.d^{2}.\log\left(\frac{4K}{\delta}\right)\left(\frac{\epsilon^{-4}}{\sigma_{k}^{-4}}\right)\right),$ (29)
we ensure that $A(k)\leq\frac{\epsilon^{2}}{\sigma_{k}^{2}}=\frac{\epsilon^{2}}{1-e^{-2(T-t_{k})}}$ for all $k\in\{0,\dots,K-1\}$. Summing over all time steps, we obtain with probability at least $1-\delta$
$\sum_{k=0}^{K}A(k)(t_{k+1}-t_{k})$ $\leq\sum_{k=0}^{K}\frac{\epsilon^{2}}{1-e^{-2(T-t_{k})}}\leq\int_{0}^{T}\frac{\epsilon^{2}}{1-e^{-2(T-t)}}dt\leq\epsilon^{2}T.$ (30)
Note that the term $\left(\log^{2}\left(\frac{4K}{\delta}\right)\right)$ appears in the upper bound for $n_{k}$ in equation 29 since we have to take a union bound for Lemma 2 and Lemma 3 and then take a union bound over $K$ discretization steps. Substituting this bound into equation 18, and then substituting the result into equation 16, we obtain that with probability at least $1-\delta$.
$\mathrm{TV}(p_{t_{0}},\hat{p}_{t_{0}})$ $\leq\mathcal{O}(\exp^{-T})+\mathcal{O}\left(\frac{1}{\sqrt{K}}\right)+\mathcal{O}(\epsilon\cdot\sqrt{T})+\epsilon_{approx}$ (31)
Finally, by choosing $T=\mathcal{O}\left(\log\left(\frac{1}{\epsilon}\right)\right)$, and $K=\tilde{\mathcal{O}}(\epsilon^{-2})$, we conclude that with probability at least $1-\delta$
$\mathrm{TV}(p_{t_{0}},\hat{p}_{t_{0}})\leq{\mathcal{O}}(\epsilon)+\epsilon_{approx},$ (32)
@@
In summary, our work provides a principled decomposition of the errors in score-based generative models, highlighting how each component contributes to the overall sample complexity. This leads to the first finite sample complexity bound of $\widetilde{\mathcal{O}}(\epsilon^{-4})$ for diffusion models without assuming access to the empirical minimizer of the score estimation function.